OpenStudy (kgrendel0324):

Presley is organizing the school book sale. Each stall will have 40 books on the shelf and 2 boxes of books with b books inside. Presley also has 200 books in inventory to refill the shelves as books sell. This situation can be written as the following expression: 40s + (2b)s + 200 How many books should Presley have ready if she plans to open 8 stalls and have 12 books in each box?

OpenStudy (yanasidlinskiy):

T=40S+(2b)S+200 T=S(40+2b)+200 T=8(40+2(12))+200 T=8(40+24)+200 T=8(64)+200 T=512+200=712 :)

OpenStudy (mathstudent55):

Just substitute s with 8 and b with 12, and evaluate the expression. 40s + (2b)s + 200 = = 40(8) + 2(12)(8) + 200 = 320 + 192 + 200 = 712